OpenStudy (anonymous):

what is the value of the function y=3/2t-5/2 when t = 3 I will medal/fan

OpenStudy (ibbutibbu.):

\[y=\frac{ 3 }{ 2 }t-\frac{ 5 }{ 2 }\]

OpenStudy (ibbutibbu.):

Is this correct?

OpenStudy (ibbutibbu.):

If so, then just substitute the value 3 where there is a t

OpenStudy (ibbutibbu.):

\[y=\frac{ 3 }{ 2 }(3)-\frac{ 5 }{ 2 }\]

OpenStudy (ibbutibbu.):

What is \[\frac{ 3 }{ 2 }\] times 3?

OpenStudy (anonymous):

So its like \[\frac{ 3 }{ 2 }\left( 3 \right)-\frac{ 5 }{ 2 }\]

OpenStudy (ibbutibbu.):

Precisely.

OpenStudy (anonymous):

wait.. \[\left[\begin{matrix}3 & 3 \\ 2 & 0\end{matrix}\right]\]

OpenStudy (anonymous):

I'm not good at math

OpenStudy (anonymous):

9 over 0?

OpenStudy (ibbutibbu.):

No, you cant divide anything by Zero

OpenStudy (anonymous):

... so its 9 over 2?

OpenStudy (ibbutibbu.):

Yes, I think so...

OpenStudy (ibbutibbu.):

then, you add 9/2 and 5/2 to get...?

OpenStudy (anonymous):

So that means \[\frac{ 9 }{ 2 }-\frac{ 5 }{ 2 }=\frac{ 4 }{ 2 }\ 4\div2=2?\]

OpenStudy (ibbutibbu.):

Not exactly....

OpenStudy (anonymous):

Ugh

OpenStudy (ibbutibbu.):

Oh no wait

OpenStudy (ibbutibbu.):

You are right, it IS 2

OpenStudy (anonymous):

:D

OpenStudy (ibbutibbu.):

Nice job!

OpenStudy (anonymous):

Thanks ^-^ Can you help me with two more? Well if you want to >.>'

OpenStudy (anonymous):

I'm just askin bruh... not pullin chains

OpenStudy (ibbutibbu.):

Sure! But I only have 2 minutes

OpenStudy (anonymous):

darn it... okay brb

OpenStudy (ibbutibbu.):

If it;s Quick

OpenStudy (anonymous):

Evaluate the function y = 17 – x when x = –6 A. 102 B. 11 C. –23 D. 23

OpenStudy (ibbutibbu.):

Uhhhh, what?

OpenStudy (ibbutibbu.):

Are those question marks supposed to be there, or di just ignore them

OpenStudy (anonymous):

Just ignore em Find the function when p=6 p=0.008n

OpenStudy (anonymous):

Here are the choices A. 0.048 B. 0.0042 C. 0.48 D. 6.008

OpenStudy (anonymous):

Oops i meant n=6

OpenStudy (ibbutibbu.):

A

OpenStudy (ibbutibbu.):

Gotta go now, Bye

OpenStudy (anonymous):

Bye thanks for the help